# A Survey of Progress and Challenges in Aligning LLMs with Human Intentions

### v2

@@ 3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments Cheang et al. (2023); Xu et al. (2024). From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 7 Q.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 7 Q.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL Jang et al. (2022) ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness Liska et al. (2022). RKC-LLMs Wang et al. (2023) allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information. BIPIA Yi et al. (2025) used adversarial training to combat the effects of information pollution and improve model robustness. CAR Weller et al. (2024) achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system.
@@ 3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks Kim et al. (2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 8 Q.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve this problem, researchers have started to parse the semantic information expressed by users through clue engineering. Folkscope Yu et al. (2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko Lu et al. (2024) introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. ATC Deng et al. (2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
@@ 3.4 Intention Clarification Failure
To solve above problems, researchers have started to try to construct multi-domain datasets Chen et al. (2023) containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 Guo et al. (2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 Muennighoff et al. (2025) uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds "wait" to guide in-depth reasoning and optimize the quality of the answer. MoChat Mo et al. (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
Cat. Benchmark Year Lang. Num. Type Description
Instrution Understanding BotChat Duan et al. (2024) 2023 En&Zh 7658 M Multi-round dialogue eval. via simulated data
MINT Wang et al. (2024) 2023 En 29,307 M Tool use and feedback in multi-turn dialogue
MT-BENCH-101 Bai et al. (2024) 2024 En 1,388 M Multi-turn dialogue ability
$\infty${B}ench Zhang et al. (2024c) 2024 En&Zh 100k S Long-context handling
L-Eval An et al. (2024) 2024 En 2,000 S Evaluation of Long-Context Language Models
LongICLBench Li et al. (2025) 2025 En 2,100k S Long In-Context Learning
Intention Reasoning BIPIA Yi et al. (2025) 2023 En 712.5K S Vulnerability to hint injection
Miko Lu et al. (2024) 2024 En 10k S Multimodal social intent understanding
CONTRADOC Li et al. (2024a) 2023 En 891 S Self-contradiction in long docs
CDCONV Zheng et al. (2022) 2022 Zh 12K M Contradiction in Chinese dialogues
Relibale Dialog Generate Open-LLM-Leaderboard Ye et al. (2024) 2024 En 10K S Uncertainty in generation
ETHICS Hendrycks et al. (2021) 2020 En 130K S Moral reasoning
FACTOR Muhlgay et al. (2024) 2023 En 300 S Factuality in generated text
@@ 4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al. (2019). If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 10 Q.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 10 Q.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs.
Fine-tuning LLMs.
To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models Sensoy et al. (2018); Amini et al. (2020). LUQ Zhang et al. (2024a) is a novel sampling-based uncertainty quantification method specifically designed for long texts. ConformalFactuality Mohri and Hashimoto (2024) defines the associated uncertainties for each possible output.
External Tools.
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al. (2024). ConfidenceElicitation Xiong et al. (2024) is a new uncertainty measurement tool for large model outputs. CalibrateMath Lin et al. assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
@@ 4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al. (2021). This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 11 Q.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al. (2024); Dutta et al. (2024). 2) Conflicts with Moral/Ethical Standards (Figure 11 Q.1). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al. (2024). To tackle the above concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
Pretraining Data Cleaning and Curation.
@@
To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes Bender et al. (2021), such as filtering out toxic language, hate speech, and harmful stereotypes. Tools like word embedding debiasing methods Rakshit et al. can help identify and remove toxic and biased content.
Reinforcement Learning-based Alignment.
@@
To further align LLMs with human and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al. (2022), DPO Zeng et al. (2024), and GRPO Shao et al. (2024), are very essential. As an extension of RLHF, RLAIF Lee et al. (2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
In-context Alignment.
@@
It leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. Lin et al. (2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. Huang et al. (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
5 Case Analysis
To demonstrate the challenges of existing LLMs in above stages, we provide cases analysis in Appendix B, and make statistical comparisons in Table 2. The results show that models with reasoning capabilities (such as GPT-o3*, Deepseek-R1*) perform better in command understanding tasks such as long text understanding and multi-round dialogue, but have not completely solved all problems. In the intention reasoning task, except for "Inconsistent information reasoning", the models generally performed poorly, and only a few models partially passed the case. In terms of reliable dialogue generation, all models performed poorly in response stability, and only GPT-o3 and Deepseek-R1 performed well in the alignment test. Overall, while reasoning abilities contribute to improved task performance, significant limitations remain across all three stages.
6 Benchmark
This section covers benchmarks for LLMs in above three stage (Table 1), more details see Appendix A.
6.1 Benchmarking Instruction Understanding
Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues. LLM capabilities in multi-turn dialogue and long-context processing have been explored through various benchmarks. BotChat Duan et al. (2024) evaluates dialogue generation, showing GPT-4’s strengths but noting instruction compliance and length limitations in other models. MINT Wang et al. (2024) highlights limited progress in tool use and feedback for complex tasks. MT-Bench-101 Bai et al. (2024) identifies challenges in enhancing long-term interaction skills. For long-context tasks, performance drops in ultra-long texts (Zhang et al. Zhang et al. (2024c)), L-Eval An et al. (2024) emphasizes length-instruction-enhanced metrics, and LongICLBench Li et al. (2025) reveals difficulties in reasoning across multiple pieces of information.
@@ 6.1 Benchmarking Instruction Understanding
6.2 Benchmarking LLM Reasoning
LLM intention reasoning involves inferring user intentions by interpreting both explicit and implicit language cues. The existing benchmarks comprehensively evaluate the multifaceted reasoning capabilities of LLMs, encompassing vulnerability to indirect hint injection attacks (BIPIA Yi et al. (2025)), advantages in multimodal intention understanding (Miko Lu et al. (2024)), analysis of self-contradictions in long documents (CONTRADOC Li et al. (2024a)), and contradiction detection in Chinese dialogues (CDCONV Zheng et al. (2022)). Collectively, these benchmarks highlight both the challenges and advancements of LLMs in complex reasoning.
6.3 Benchmarking LLM Generation
LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses. Open-LLM-Leaderboard Ye et al. (2024), ETHICS Hendrycks et al. (2021) and FACTOR Muhlgay et al. (2024) all focus on the reliability evaluation of content generated by large models. Open-LLM-Leaderboard finds that large-scale models have higher uncertainty, and fine-tuned models have higher accuracy but greater uncertainty. ETHICS focuses on the ethical value alignment of generated content. FACTOR evaluates factuality through scalable methods to ensure that diverse and rare facts are covered.
7 Future Directions
Automated Annotation Framework.
GraphRAG.
Balancing Safety and Performance.
8 Conclusion
Limitations
We have made significant efforts to ensure the quality of this study, but certain limitations may still exist. First, due to constraints on the length of the discussion, this article primarily focuses on analyzing examples of human interaction with large models, which inevitably limits the depth of methodological details we can explore. Second, while our meta-analysis draws extensively from prominent academic platforms such as the ACL conference series, ICLR, ICML, and the arXiv preprint repository, it is possible that some valuable insights from other sources have not been included. It is worth noting that several open scientific questions in this field remain unresolved, and the academic community has yet to reach a consensus on these issues. To address these limitations, we plan to establish a long-term tracking mechanism to monitor new developments in the field. This will allow us to incorporate emerging theoretical advancements and dynamically refine or expand upon the perspectives presented in this study.
Appendix A Benchmark Details
A.1 Instruction Understanding Benchmark
BotChat Duan et al. (2024) specifically addresses the evaluation of LLMs’ ability to emulate human-like, multi-turn conversations using an LLM-centric approach. This benchmark assesses utterance generation, evaluation protocols, and judgment. Findings suggest that models like GPT-4 demonstrate exceptional performance as both generators and judges, producing human-indistinguishable dialogues and showing high alignment with human evaluation standards. Conversely, other LLMs face challenges in generating quality multi-turn dialogues due to issues like inadequate instruction-following and a tendency towards prolixity, particularly in generating extensive dialogues.
MINT Wang et al. (2024) introduces a benchmark that evaluates LLMs’ capacity to solve complex tasks through multi-turn interactions, emphasizing the use of external tools and leveraging natural language feedback. MINT provides a reproducible evaluation framework where LLMs can access tools and receive simulated user feedback. Analysis using MINT reveals that LLMs generally benefit from tools and language feedback. Interestingly, studies with MINT suggest that techniques like supervised instruction-finetuning (SIFT) and reinforcement learning from human feedback (RLHF) might not consistently improve multi-turn capabilities, indicating the need for further research in training methodologies for conversational settings.
MT-BENCH-101 Bai et al. (2024) offers a structured approach with its three-layer hierarchical capability taxonomy and a dataset of 1,388 dialogue pairs across 13 tasks. This allows for a fine-grained evaluation of LLMs’ multi-turn dialogue skills. The findings from studies utilizing MT-BENCH-101 indicate that commonly employed techniques yield limited improvements in multi-turn performance, suggesting a need for more effective methods to enhance conversational abilities over extended interactions.
Zhang et al. Zhang et al. (2024c) focus on the challenge of long-context processing, introducing a benchmark of 12 tasks, each averaging over 100K tokens, to assess LLMs’ long-context processing abilities. Results indicate a marked performance drop with longer contexts, highlighting the need for further improvement.
L-Eval An et al. (2024) contributes to standardizing the evaluation of Long-Context Language Models (LCLMs) by providing a new evaluation suite with diverse tasks, document lengths (3k to 200k tokens), and a large number of human-labeled query-response pairs. It also investigates the effectiveness of evaluation metrics, suggesting that Length-instruction-enhanced (LIE) evaluation and LLM judges correlate better with human judgments. Comprehensive studies using L-Eval offer insights into the performance of both commercial and open-source LCLMs.
LongICLBench Li et al. (2025) focuses on long in-context learning in extreme-label classification, with input lengths ranging from 2K to 50K tokens and a large number of classes. Evaluation on this benchmark showed that while LLMs perform well on less challenging tasks, they struggle with more complex ones involving a large label space and longer contexts, revealing a bias towards later parts of the input and difficulty in reasoning over multiple pieces of information.
A.2 LLM Intention Reasoning Benchmark
BIPIA Yi et al. (2025) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities.
Miko Lu et al. (2024) assesses multimodal models in understanding social media user intentions. The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4. Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy.
CONTRADOC Li et al. (2024a) is the first dataset for analyzing self-contradictions in long documents. Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning.
CDCONV Zheng et al. (2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds. The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
A.3 LLM Reliable Generation Benchmark
Open-LLM-Leaderboard Ye et al. (2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue. Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy.
ETHICS Hendrycks et al. (2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being. The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain.
FACTOR Muhlgay et al. (2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered. Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
Appendix B Challenges Cases in Human-LLMs Aligement
@@
To fully understand the limitations of large language models (LLMs) in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios. We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing large language models (LLMs) face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
@@
B.1 Case of Instruction Understanding
Although large language models (LLMs) perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.

### v3

@@ 3.2 Misinformation Reasoning
Erroneous instructions mislead model outputs more severely than inconsistent ones, as they lack obvious contradictions, requiring the model to comprehend, reason, compare input knowledge with its parameterized knowledge, and make objective judgments [102, 103]. From the input perspective, this paper classifies the sources of erroneous information into two categories as follows: 1) Temporal Alignment Failure (Figure 7 Q.1). arises when the knowledge provided by the user and the model is temporally misaligned due to updates occurring at different times, leading to inconsistent responses. Such discrepancies typically originate during the training process. 2) Information Contamination (Figure 7 Q.2). refers to the degradation of model quality caused by the intentional distortion of input data.
To solve the above problems, existing methods mainly focus on improving model susceptibility in the face of internal and external knowledge conflicts through targeted fine-tuning and processing. CKL [40] ensures that the model’s knowledge is updated in a timely manner through an online approach, although this approach is slightly weaker than re-training in terms of effectiveness [41]. RKC-LLMs [46] allows a large model to recognize knowledge conflicts by means of instructional fine-tuning that identifying specific passages of conflicting information. BIPIA [42] used adversarial training to combat the effects of information pollution and improve model robustness. CAR [47] achieved nearly 20% improvement by discriminating external knowledge that may not be contaminated in the RAG system. See Figure 3 for more methods.
@@ 3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks [50], and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 8 Q.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve this problem, researchers have started to parse the semantic information expressed by users through clue engineering. Folkscope [48] proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko [49] introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. Zhang et al. [53] have employed behavioral cloning by using demonstration data from a strong model to train a weaker model so that the weaker model can perform better on a similar task. ATC [54] utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain. See Figure 3 for more methods.
@@ 3.4 Intention Clarification Failure
To solve above problems, researchers have started to try to construct multi-domain datasets [57] containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 [55] enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 [56] uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer. SoulChat [57] fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations. MoChat [58] constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization. LARA [68] combines a fine-tuned smaller model with retrieval enhancement mechanisms and integrates it into the architecture of LLM. See Figure 3 for more methods.
@@ 4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want [104]. If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 10 Q.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 10 Q.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs.
Fine-tuning LLMs
To make LLMs more accurate in estimating uncertainty, existing methods fine-tune models [69, 70]. LUQ [80] is a novel sampling-based uncertainty quantification method specifically designed for long texts. UaIT [79] using semantic entropy to assess output uncertainty. ConformalFactuality [71] defines the associated uncertainties for each possible output. See Figure 3 for more methods.
External Tools
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools [81]. ConfidenceElicitation [105] is a new uncertainty measurement tool for large model outputs. CalibrateMath [82] assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
@@ 4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content [84]. This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 11 Q.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics [106, 107]. 2) Conflicts with Moral/Ethical Standards (Figure 11 Q.1). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities [108, 109]. To tackle the above concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
Pretraining Data Cleaning and Curation
@@
To minimize the risks associated with harmful or inappropriate content, LLM training datasets should undergo rigorous cleaning processes [84], such as filtering out toxic language, hate speech, and harmful stereotypes. Tools like the Perspective API [83] and word embedding debiasing methods [85] can help identify and remove toxic and biased content. Tools like word embedding debiasing methods [85] can help identify and remove toxic and biased content.
Reinforcement Learning-based Alignment
@@
To better align LLMs with human values and societal norms, reinforcement learning approaches are widely adopted, including RLHF and its advanced variants such as PPO [3], DPO [86], and GRPO [87]. Extending RLHF, RLAIF [88] leverages AI systems to assist in the feedback process, improving the scalability of evaluation and fine-tuning while enabling continuous monitoring and refinement of model behavior [88].
In-context Alignment
@@
It leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. [89] demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. [90] explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities. Liu et al. [91] introduced PICA, which uses a two-stage approach to improve alignment efficiency.
@@
5 Challenges Cases in Human-LLMs Aligement
@@
To fully understand the limitations of LLMs in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios. We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing LLMs face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
@@
5.1 Case of Instruction Understanding
Although LLMs perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.